Write 780 as a product of its primes. Use exponents in your notation
Answer by Theo(13342)

780 = 2 * 2 * 5 * 3 * 13 = 2^2 * 5 * 3 * 13

solution is 2^2 * 5 * 3 * 13

list of prime numbers up to 100.
2,3,5,7,11,13,17,19,23,29,31,37,41,43,47,53,59,61,67,71,73,79,83,89,97